# Geometry-aware 4D Video Generation for Robot Manipulation

### v1

@@ Abstract
Understanding and predicting the dynamics of the physical world can enhance a robot’s ability to plan and interact effectively in complex environments. While recent video generation models have shown strong potential in modeling dynamic scenes, generating videos that are both temporally coherent and geometrically consistent across camera views remains a significant challenge. To address this, we propose a 4D video generation model that enforces multi-view 3D consistency of videos by supervising the model with cross-view pointmap alignment during training. This geometric supervision enables the model to learn a shared 3D representation of the scene, allowing it to predict future video sequences from novel viewpoints based solely on the given RGB-D observations, without requiring camera poses as inputs. Compared to existing baselines, our method produces more visually stable and spatially aligned predictions across multiple simulated and real-world robotic datasets. We further show that the predicted 4D videos can be used to recover robot end-effector trajectories using an off-the-shelf 6DoF pose tracker, supporting robust robot manipulation and generalization to novel camera viewpoints.
@@ 1 Introduction
Understanding how the visual world changes with interactions is a key capability for intelligent robotic systems. For robot manipulation tasks, the robots need to anticipate how the environment changes by taking into account object motions or occlusions upon interactions over extended time horizons. Recent advancements in video generation models offer a powerful paradigm for learning such dynamic visual models: by forecasting future observations, robots can simulate possible outcomes, plan actions, and adapt to new environments.
However, it remains a challenge to generate realistic and physically plausible videos which smooth and precise robot policies can be extracted from. There are two challenges: temporal coherence, ensuring smooth, causally consistent motion over time; and 3D consistency, preserving object geometry and spatial correspondences across different viewpoints. Most existing video generation models capture one at the expense of the other. Pixel-based models [1, 2] trained on RGB videos often excel at short-term motion but lack an understanding of 3D structures, which leads to artifacts like flickering, deformation, or object disappearance. In contrast, 3D-aware approaches enforce geometric constraints but are limited to simple, static backgrounds and struggle to scale to realistic, multi-object manipulation scenarios [3, 4, 5].
In this work, we present a video generation framework that bridges this gap by unifying strong temporal modeling with robust 3D geometric consistency. Our method produces 4D videos that can be rendered into RGB-D sequences that are both coherent over time and spatially consistent across camera views. To achieve this, we introduce a geometry-consistent supervision mechanism inspired by DUSt3R [6] and adapt it for the video generation task. Specifically, the model is trained to predict a pair of 3D pointmap sequences: one for a reference view and one for a second view projected into the reference view camera coordinate frame. By minimizing the difference between reference and projected 3D points over time, the model learns a shared scene representation across views. This enables robust generalization to novel viewpoints at inference, which is particularly useful in robotic applications where even small camera view shifts can push visuomotor policies out of distribution and lead to failures.
@@
We further demonstrate that the predicted multi-view RGB-D videos can be directly used to extract robot end effector trajectories using an off-the-shelf 6DoF pose tracker, such as FoundationPose [7]. We evaluate this approach on three simulated robot manipulation tasks where the camera views are unseen during training, achieving good success rates on all tasks and outperforms the baseline method. Additionally, our generation model with geometry-consistent supervision allows any novel view to be projected into the known reference frame without additional camera calibration, enabling flexible camera placement and simplifying robot deployment.
In summary, we propose a 4D video generation framework that achieves both 3D geometric consistency and temporal coherence. To achieve this, we first introduce a geometry-consistent supervision mechanism that enforces cross-view alignment of generated videos over time. Second, we develop a benchmark for video generation in robotic manipulation, comprising both simulation and real-world tasks. Each task is recorded from diverse camera viewpoints, enabling comprehensive evaluation of 4D generation quality and generalization to unseen views. Finally, we demonstrate that the generated 4D videos can be directly used to extract robot trajectories using an off-the-shelf 6DoF pose tracker, enabling reliable manipulation under novel viewpoints without additional camera calibration.
@@ 2 Related Work
Video Generation has been a long-standing task in computer vision. Early works use recurrent networks [8, 9] or generative adversarial networks [10] to learn temporal dynamics from video data. With the recent success of image diffusion models, many works have extended diffusion models for the video prediction task by adding additional temporal layers [1, 2] and using latent diffusion techniques [11, 12, 13, 14]. In addition to RGB videos, video diffusion models have also been shown to generate other modalities such as high-fidelity depth videos [15, 16, 17] and pointmaps [18]. In this work, we adapt a latent video diffusion model (SVD) to jointly learn from RGB and depth videos with improved temporal and spatial consistency.
@@
Multi-View and 4D Video Generation. Recent advances in camera-conditioned video generation improve spatial consistency by associating camera poses with multi-view videos [19, 20, 21, 22]. However, 4D video generation requires joint reasoning about both object geometry across views and motion over time. Prior 4D generation works separately optimize for temporal consistency with a video model and spatial consistency with a novel view synthesis model [23, 24, 25]. In this work, we present a 4D video generation framework unifies the two objectives. In particular, our method introduces a geometry-consistent supervision mechanism through cross-view pointmap alignment, inspired by DUSt3R [6], on top of the standard video diffusion process. This approach enforces spatial consistency across camera views while ensuring temporal coherence. Our goal aligns with recent 4D generation works that jointly optimize spatial and temporal consistency [3, 4], but while they primarily focus on single-object videos with white backgrounds, we target multi-object, dynamic robot manipulation scenes.
@@
Generative Models for Robot Planning. With the recent success of video generation models and their strong generalization capabilities across diverse visual scenes, many works have explored their potential as dynamics models for robotic tasks. Specifially, robot actions can be extracted from predicted future frames using a learned inverse dynamics model [26, 27, 28], a behavior cloning policy conditioned on generated outputs [29, 30, 31], or an RGB-based pose tracking model [32]. To more tightly couple future state prediction with action inference, recent methods have proposed unified models that simultaneously predict both future video frames and robot actions [33, 34]. Yet it remains a challenge to generate spatially consistent predictions across views and capture accurate 3D geometry needed for precise robot manipulation. To bridge this gap, we propose a model that learns spatial correspondences across camera views, enabling better RGBD video generation quality on novel views and improved end effector pose tracking accuracy for better robot task performance.
@@ 3.1 Diffusion-Based Video Generation
We adopt the Stable Video Diffusion [11] framework which has demonstrated strong performance in generating short, temporally coherent video sequences. It first projects historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ into a latent space using a pretrained Variational Autoencoder (VAE) [35] encoder. A diffusion model, implemented as a U-Net with an encoder-decoder structure, then predicts future latent representations $\{\mathbf{z}_{t+1},\cdots,\mathbf{z}_{t+h}\}$, which are decoded back into RGB frames $\{\mathbf{O}_{t+1},\cdots,\mathbf{O}_{t+h}\}$ using VAE decoder. The diffusion model $f_{\theta}$ is trained using an alternative of the standard DDPM [36] method, which directly predicts the original clean data from the noisy input at each diffusion step. The training objective for predicting a future latent $z_{t^{\prime}}$ at timestep $t^{\prime}$ is to minimize:
@@ 3.2 Geometry-Consistent Supervision
To enforce 3D consistency across views, we adopt the cross-view pointmap supervision strategy from DUSt3R [6], adapted to the video generation setting. As shown in Figure 2, given the history pointmaps $\{\mathbf{X}_{t-h+1}^{n},\cdots,\mathbf{X}_{t}^{n}\}$ from camera view $v_{n}$, we first encode them using a Pointmap VAE, which is initialized from the pretrained RGB VAE from SVD [11] and fine-tuned on pointmap data. This produces the latent representation $\{\mathbf{z}_{t+1}^{n},\cdots,\mathbf{z}_{t+h}^{n}\}$. We then apply the same latent diffusion method used for RGB video prediction to forecast future pointmaps in the latent space. The predicted latents are subsequently decoded by the Pointmap VAE decoder to obtain future pointmaps $\{\mathbf{X}_{t+1}^{n},\ldots,\mathbf{X}_{t+h}^{n}\}$.
@@
where $\mathbf{z}_{t^{\prime}}^{n}(k)$ denotes the noised latent of the pointmap from view $v_{n}$ at diffusion step $k$, and $\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(k)$ is the noised latent for the pointmap from view $v_{m}$ projected into the coordinate frame of view $v_{n}$. Similar to Equation 1, $\epsilon_{t^{\prime}}^{n}$ and $\epsilon_{t^{\prime}}^{m}$ are Gaussian noise added to the pointmap latents from view $v_{n}$ and $v_{m}$, respectively. The conditioning latents $c^{n}$ and $c^{m}$ are derived from the historical pointmaps of their respective views. See Appendix for model architecture details.
@@ 3.4 Robot Pose Estimation from 4D videos
We leverage the predicted 4D video to extract robot trajectories using an off-the-shelf 6DoF pose tracking model, FoundationPose [7]. The model takes as input RGB-D frames from a single view, a binary mask of the target object in the initial frame (generated using SAM2 [37]), camera intrinsics, and the gripper CAD model. At each pose estimation timestep, the model outputs the estimated pose of the CAD model, $\mathbf{T}_{t}\in\text{SE}(3)$, along with a confidence score for the prediction, and tracks the object for subsequent frames.
Pose estimation is run independently for both views, and the result with the highest confidence score is selected. Since all outputs are aligned to the coordinate frame of the first camera view, only the first camera’s extrinsic must be calibrated. This eliminates the need for calibrating additional cameras at test time and enables flexible multi-view configurations.
To infer the gripper open/close state, we segment the left and right gripper fingers and project their pixels into 3D space based on the predicted RGB-D sequences. The distance between the centroids of the two finger point clouds is measured: if it falls below a threshold $\delta$, the gripper is considered closed; otherwise, it is considered open. The recovered trajectories are directly used to control the robot to execute downstream tasks.
@@ 4.1 Tasks
We evaluate the 4D video generation results on three simulated tasks and one real-world task. The tasks are drawn from the LBM environment [38], a physics-based simulator built on Drake [39] that provides realistic rendering and demonstrations for both single-arm and bimanual robot manipulation.
Simulation Task: The simulation tasks are illustrated in Figure 3: StoreCerealBoxUnderShelf, PutSpatulaOnTable, and PlaceAppleFromBowlIntoBin. In StoreCerealBoxUnderShelf, a single robot arm picks up a cereal box from the top of a shelf and inserts it into the shelf below. Occlusions occur during insertion, especially from certain camera viewpoints, making multi-view predictions essential. Additionally, the pick-and-insert action requires spatial understanding and precision. In PutSpatulaOnTable, the robot arm retrieves a spatula from a utensil crock and places it on the left side of the table. This task requires precise manipulation to successfully grasp the narrow object. In PlaceAppleFromBowlIntoBin, one arm picks up an apple from a bowl on the left side of the table and places it on a shelf; a second arm then picks up the apple and deposits it into a bin on the right side. This is a long-horizon, bimanual task that tests the model’s ability to predict both temporally and spatially consistent trajectories.
The dataset consists of 25 demonstrations for the StoreCerealBoxUnderShelf task, 22 demonstrations for the PutSpatulaOnTable task, and 20 demonstrations for the PlaceAppleFromBowlIntoBin task, with different initial object configurations. Each demonstration includes RGB-D observations from 16 different camera poses. We randomly sample 12 views for training and reserve the remaining 4 for testing. All camera poses are sampled from the upper hemisphere positioned above the workstation. More details and visualization can be found in Section A.2.
Real-world Task: The real-world task mirrors the PutSpatulaOnTable task from simulation. We collected 5 demonstration videos using two FRAMOS D415e cameras positioned at different viewpoints. Each camera records synchronized RGB-D observations of the manipulation sequence. This setup allows us to evaluate the model’s ability to generate real-world robot videos.
@@
Task 4: Real World PutSpatulaOnTable
OURS 0.58 232.31 239.06 0.06 0.08 0.97 0.95
OURS w/o MV attn 0.34 266.94 254.34 0.11 0.11 0.89 0.89
4D Gaussian [40] 0.00 2002.48 2708.86 0.30 1.75 0.82 0.26
SVD [11] 293.07 319.06
SVD w/ MV attn 288.12 257.90
Table 1: Multi-view 4D Video Generation Results under Novel Camera Views. We compare our method with baselines in terms of cross-view consistency, RGB video generation quality, and depth generation quality. Our method consistently enables high-quality video and depth generation while maintaining strong cross-view consistency on both simulated and real-world datasets.
Figure 4: Qualitative Results and Comparisons under Novel Camera Views. Our method generates geometrically consistent 4D videos across camera views. In contrast, baseline results often exhibit significant cross-view inconsistencies or contain noticeable artifacts in the RGB or depth predictions.
@@
Video prediction: To evaluate RGB video generation quality, we compute the commonly used Fréchet Video Distance [41] (FVD) between the generated video and ground-truth video. FVD-$n$ evaluates the prediction from the reference view $v_{n}$, and FVD-$m$ evaluates the predcition from view $v_{m}$.
@@
Cross-View 3D Consistency: To evaluate the 3D consistency of the generated pointmaps across views, we compute the mean Intersection-over-Union (mIoU) on object masks. We use SAM2 [37] to track the robot gripper and obtain binary masks from the generated videos for both views. For each frame, we lift the gripper mask in view $v_{n}$ to 3D space and then re-project it to view $v_{m}$. We then compute the IoU between the bounding boxes of the projected and original gripper masks. The mIoU is averaged over all time steps, and higher values indicating stronger 3D alignment across views.
Baselines. We compare our method with prior 4D generation approaches and variants of our model to evaluate generation quality and multi-view consistency. All models are trained on the same multi-view RGB-D video dataset and tested on novel viewpoints not observed during training.
@@
  • 4D Gaussian [40]: A baseline method that predicts one single-view RGB video using a finetuned SVD model [11] on our dataset, and then use a 4D Gaussian method, Shape of Motion [40], to reconstruct a dynamic 4D scene from the video.
  • SVD [11]: Stable Video Diffusion is a state-of-the-art video generation model. We finetune it on our dataset to predict stereo RGB video sequences.
@@
Results on Simulation Tasks. Results are reported in Table 1, and Figure 4 visualizes the predictions produced by different methods. All results are evaluated on novel views not seen during training. Our method OURS consistently achieves the best or highly competitive results across all tasks. For multi-view RGB video generation, it produces lower FVD scores than all baselines in most cases, indicating high temporal coherence and visual fidelity. Additionally, our method achieves the best depth prediction and cross-view consistency scores, demonstrating that the geometry-consistent supervision effectively guides the model to generate spatially aligned and geometrically coherent videos.
@@
We show that multi-view cross-attention is a crucial design choice for helping the model learn 3D geometric correspondences across camera views. The variant without multi-view cross-attention OURS w/o MV attn exhibits significantly lower 3D consistency, as measured by the mIoU metric. The video and depth generation quality also degrades in both views. In particular, the model performs poorly on the AbsRel-$m$ and $\delta_{1}$-$m$ metrics, indicating that it fails to learn the transformation needed to generate pointmaps in view $v_{n}$’s camera coordinate frame from view $v_{m}$ without the support of cross-attention layers. In Figure 4, the gripper pose of OURS w/o MV attn is inconsistent across the two RGB views, and the projected gripper mask significantly misaligns with the actual gripper mask, as shown in the last column.
@@
4D Gaussian performs worse in video and depth generation quality, as well as cross-view consistency, as shown in Table 1. In Figure 4, the generated RGB frames appear blurry from unseen viewpoints, and the predicted depth is inaccurate. This is because the model is optimized on one single generated RGB video with fixed camera view, making it difficult to synthesize novel views.
@@
For the other baselines, SVD and SVD w/ MV attn, depth is not predicted and no geometric supervision is applied. While multi-view cross-attention is added to SVD w/ MV attn to enable information transfer between camera views, it operates over RGB features rather than pointmaps—which naturally encode 3D structure, as used in our method. As a result, the generated videos lack 3D geometric consistency across views. As shown in Figure 4, SVD w/ MV attn produces a noticeable gripper mismatch between the two views and the RGB generation quality is significantly worse than our method.
@@
Results on Real-world Task. We also evaluate our method on the real-world PutSpatulaOnTable task, as shown in Table 1. We use the checkpoint trained on the same task in simulation and finetune it on 5 real-world demonstration videos. In Figure 5, we show qualitative RGB-D prediction results. The model is able to generate high-fidelity future RGB-D sequences, accurately capturing both visual appearance and depth over time comparing to the baselines.
Figure 5: Real World 4D Video Generation Results on PutSpatulaOnTable. Our model predicts high-fidelity RGB-D sequences that capture the robot gripper motions. In this particular sequence, the model correctly predicts how the robot reaches the spatula, grasps it, and lifts it up from the utensil crock.

### v2

@@ Abstract
Understanding and predicting dynamics of the physical world can enhance a robot’s ability to plan and interact effectively in complex environments. While recent video generation models have shown strong potential in modeling dynamic scenes, generating videos that are both temporally coherent and geometrically consistent across camera views remains a significant challenge. To address this, we propose a 4D video generation model that enforces multi-view 3D consistency of generated videos by supervising the model with cross-view pointmap alignment during training. Through this geometric supervision, the model learns a shared 3D scene representation, enabling it to generate spatio-temporally aligned future video sequences from novel viewpoints given a single RGB-D image per view, and without relying on camera poses as input. Compared to existing baselines, our method produces more visually stable and spatially aligned predictions across multiple simulated and real-world robotic datasets. We further show that the predicted 4D videos can be used to recover robot end-effector trajectories using an off-the-shelf 6DoF pose tracker, yielding robot manipulation policies that generalize well to novel camera viewpoints.
@@ 1 Introduction
Understanding how the visual world changes with interactions is a key capability for intelligent robotic systems. For robot manipulation tasks, the robots need to anticipate how the environment changes by taking into account object motions or occlusions upon interactions over extended time horizons. Recent advancements in video generation models offer a powerful paradigm for learning such dynamic visual models: by forecasting future observations, robots can simulate possible outcomes, plan actions, and adapt to new scenarios.
However, it remains a challenge to generate realistic and physically plausible videos from which smooth and precise robot policies can be extracted from. There are two challenges: temporal coherence, ensuring smooth, causally consistent motion over time; and 3D consistency, preserving object geometry and spatial correspondences across different viewpoints. Most existing video generation models capture one at the expense of the other. Pixel-based models (Yan et al., 2021; Ho et al., 2022) trained on RGB videos often excel at short-term motion but lack an understanding of 3D structures, which leads to artifacts like flickering, deformation, or object disappearance. In contrast, 3D-aware approaches enforce geometric constraints but are limited to simple, static backgrounds and struggle to scale to realistic, multi-object manipulation scenarios (Xie et al., 2024; Li et al., 2024a; Zhang et al., 2024).
In this work, we present a video generation framework that bridges this gap by unifying strong temporal modeling with robust 3D geometric consistency. Our method produces 4D videos that can be rendered into RGB-D sequences that are both coherent over time and spatially consistent across camera views. To achieve this, we introduce a geometry-consistent supervision mechanism inspired by DUSt3R (Wang et al., 2024b) and adapt it for the video generation task. Specifically, the model is trained to predict a pair of 3D pointmap sequences: one for a reference view and one for a second view projected into the reference view camera coordinate frame. By minimizing the difference between reference and projected 3D points over time, the model learns a shared scene representation across views. This enables robust generalization to novel viewpoints at inference, which is particularly useful in robotic applications where even small camera view shifts can push visuomotor policies out of distribution and lead to failures.
@@
We further demonstrate that the predicted multi-view RGB-D videos can be directly used to extract robot end effector trajectories using an off-the-shelf 6DoF pose tracker, such as FoundationPose (Wen et al., 2024). We evaluate this approach on three simulated robot manipulation tasks where the camera views are unseen during training, achieving good success rates on all tasks and outperforms the baseline method. Additionally, our model generates future pointmaps in the reference view’s coordinate frame solely based on the initial RGB-D observation in each view, without camera poses as inputs, bypassing extrinsics calibration for novel camera poses.
In summary, we propose a 4D video generation framework that achieves both 3D geometric consistency and temporal coherence. To achieve this, we first introduce a geometry-consistent supervision mechanism that enforces cross-view alignment of generated videos over time. Second, we develop a benchmark for video generation in robotic manipulation, comprising both simulation and real-world tasks. Each task is recorded from diverse camera viewpoints, enabling comprehensive evaluation of 4D generation quality and generalization to unseen views. Finally, we demonstrate that the generated 4D videos can be directly used to extract robot trajectories using an off-the-shelf 6DoF pose tracker, enabling applications to manipulation tasks.
@@ 2 Related Work
Video Generation has been a long-standing task in computer vision. Early works use recurrent networks (Srivastava et al., 2015; Chiappa et al., 2017) or generative adversarial networks (Vondrick et al., 2016) to learn temporal dynamics from video data. With the recent success of image diffusion models, many works have extended diffusion models for the video prediction task by adding additional temporal layers (Yan et al., 2021; Ho et al., 2022) and using latent diffusion techniques (Blattmann et al., 2023; Liu et al., 2024; Zheng et al., 2024; Yang et al., 2024b). In addition to RGB videos, video diffusion models have also been shown to generate other modalities such as high-fidelity depth videos (Saxena et al., 2023; Hu et al., 2024; Shao et al., 2024) and pointmaps (Nguyen et al., 2025). In this work, we adapt a latent video diffusion model (SVD) to jointly learn from RGB and depth videos with improved temporal and spatial consistency.
@@
Multi-View and 4D Video Generation. Recent advances in camera-conditioned video generation improve spatial consistency by associating camera poses with multi-view videos (Van Hoorick et al., 2024; He et al., 2024; Kuang et al., 2024; Cheong et al., 2024). However, 4D video generation requires joint reasoning about both object geometry across views and motion over time. Prior 4D generation works separately optimize for temporal consistency with a video model and spatial consistency with a novel view synthesis model (Sun et al., 2024; Wang et al., 2024c; Yang et al., 2024a). In this work, we present a 4D video generation framework unifies the two objectives. In particular, our method introduces a geometry-consistent supervision mechanism through cross-view pointmap alignment, inspired by DUSt3R (Wang et al., 2024b), on top of the standard video diffusion process. This approach enforces spatial consistency across camera views while ensuring temporal coherence. Our goal aligns with recent 4D generation works that jointly optimize spatial and temporal consistency (Xie et al., 2024; Li et al., 2024a), but while they primarily focus on single-object videos with white backgrounds, we target multi-object, dynamic robot manipulation scenes.
@@
Generative Models for Robot Planning. With the recent success of video generation models and their strong generalization capabilities across diverse visual scenes, many works have explored their potential as dynamics models for robotic tasks (Yang et al., 2025b). Specifially, robot actions can be extracted from predicted future frames using a learned inverse dynamics model (Du et al., 2023; Tian et al., 2024; Zhen et al., 2025), a behavior cloning policy conditioned on generated outputs (Xu et al., 2024; Bharadhwaj et al., 2024; Huang et al., 2025a), or an RGB-based pose tracking model (Liang et al., 2024). To more tightly couple future state prediction with action inference, recent methods have proposed unified models that simultaneously predict both future video frames and robot actions (Zhu et al., 2025; Li et al., 2025). Yet it remains a challenge to generate spatially consistent predictions across views and capture accurate 3D geometry needed for precise robot manipulation. To bridge this gap, we propose a model that learns spatial correspondences across camera views, enabling better RGBD video generation quality on novel views and improved end effector pose tracking accuracy for better robot task performance.
@@ 3.1 Diffusion-Based Video Generation
We adopt the Stable Video Diffusion (Blattmann et al., 2023) framework which has demonstrated strong performance in generating short, temporally coherent video sequences. It first projects historical video frames $\{\mathbf{O}_{t-h+1},\cdots,\mathbf{O}_{t}\}$ into a latent space using a pretrained Variational Autoencoder (VAE) (Kingma et al., 2013) encoder. A diffusion model, implemented as a U-Net with an encoder-decoder structure, then predicts future latent representations $\{\mathbf{z}_{t+1},\cdots,\mathbf{z}_{t+h}\}$, which are decoded back into RGB frames $\{\mathbf{O}_{t+1},\cdots,\mathbf{O}_{t+h}\}$ using VAE decoder. The diffusion model $f_{\theta}$ is trained using an alternative of the standard DDPM (Ho et al., 2020) method, which directly predicts the original clean data from the noisy input at each diffusion step. The training objective for predicting a future latent $z_{t^{\prime}}$ at timestep $t^{\prime}$ is to minimize:
@@ 3.2 Geometry-Consistent Supervision
To enforce 3D consistency across views, we adopt the cross-view pointmap supervision strategy from DUSt3R (Wang et al., 2024b), adapted to the video generation setting. As shown in Figure˜2, given the history pointmaps $\{\mathbf{X}_{t-h+1}^{n},\cdots,\mathbf{X}_{t}^{n}\}$ from camera view $v_{n}$, we first encode them using a Pointmap VAE, which is initialized from the pretrained RGB VAE from SVD (Blattmann et al., 2023) and fine-tuned on pointmap data. This produces the latent representation $\{\mathbf{z}_{t+1}^{n},\cdots,\mathbf{z}_{t+h}^{n}\}$. We then apply the same latent diffusion method used for RGB video prediction to forecast future pointmaps in the latent space. The predicted latents are subsequently decoded by the Pointmap VAE decoder to obtain future pointmaps $\{\mathbf{X}_{t+1}^{n},\ldots,\mathbf{X}_{t+h}^{n}\}$.
@@
where $\mathbf{z}_{t^{\prime}}^{n}(k)$ denotes the noised latent of the pointmap from view $v_{n}$ at diffusion step $k$, and $\mathbf{z}_{t^{\prime}}^{m\rightarrow n}(k)$ is the noised latent for the pointmap from view $v_{m}$ projected into the coordinate frame of view $v_{n}$. Similar to Equation˜1, $\epsilon_{t^{\prime}}^{n}$ and $\epsilon_{t^{\prime}}^{m}$ are Gaussian noise added to the pointmap latents from view $v_{n}$ and $v_{m}$, respectively. The conditioning latents $c^{n}$ and $c^{m}$ are derived from the historical pointmaps of their respective views. See Appendix for model architecture details.
@@ 3.4 Robot Pose Estimation from 4D videos
We leverage the predicted 4D video to extract robot trajectories using an off-the-shelf 6DoF pose tracking model, FoundationPose (Wen et al., 2024). The model takes as input RGB-D frames from a single view, a binary mask of the target object in the initial frame (generated using SAM2 (Ravi et al., 2024)), camera intrinsics, and the gripper CAD model. At each pose estimation timestep, the model outputs the estimated pose of the CAD model, $\mathbf{T}_{t}\in\text{SE}(3)$, along with a confidence score for the prediction, and tracks the object for subsequent frames.
Pose estimation is performed independently on both views, and the result with the highest confidence score is selected. With the camera extrinsics of the reference view, the tracked poses are transformed into the global frame and used to command the robot.
To infer the gripper open/close state, we segment the left and right gripper fingers and project their pixels into 3D space based on the predicted RGB-D sequences. The distance between the centroids of the two finger point clouds is measured: if it falls below a threshold $\delta$ (0.10m for StoreCerealBoxUnderShelf, 0.06m for PutSpatulaOnTable, for 0.12m PutAppleFromBowlIntoBin), the gripper is considered closed; otherwise, it is considered open. The recovered trajectories are directly used to control the robot to execute downstream tasks.
@@ 4.1 Tasks
We evaluate 4D video generation results in both the LBM simulation and real world. The LBM simulation (Institute, 2025) is a physics-based simulator built on Drake (Tedrake and the Drake Development Team, 2019) that provides realistic rendering and demonstrations for robot manipulation tasks.
Simulation Task: We evaluate on three table-top manipulation tasks with both object and motion diversity (Figure˜6): StoreCerealBoxUnderShelf, PutSpatulaOnTable, and PlaceAppleFromBowlIntoBin. In StoreCerealBoxUnderShelf, a robot arm picks up a cereal box from the top of the shelf and inserts it into the shelf. This task requires multi-view perception due to occlusions. PutSpatulaOnTable involves grasping a spatula on its handle from a utensil crock and placing it on the table. PlaceAppleFromBowlIntoBin is a long-horizon bimanual task where one arm picks up an apple from a bowl and places it on the shelf, and then the other arm picks up the apple from the shelf and places it into a bin. The simulation dataset contains 25 demonstrations per task with varying initial object poses. Each demo includes RGB-D observations from 16 camera poses, yielding 400 videos in total per task. We use 12 views for training and 4 for testing, yielding a total of 300 camera views for training and 100 camera views for testing per task. All camera poses are sampled from the upper hemisphere positioned above the workstation. More details can be found in Section˜A.2.
Real-world Tasks: The dataset contains 4 tasks (Figure˜8) collected using two Franka Panda robot arms —AddOrangeSlicesToBowl, PutCupOnSaucer, TwistCapOffBottle, and PutSpatulaOnTable. Each task contains 20 robot teleoperation demonstrations. For each demonstration, we use two FRAMOS D415e cameras to capture synchronized RGB-D sequences from different viewpoints.
@@
Multi-task Real World Dataset
OURS 0.56 384.26 331.00 0.10 0.20 0.93 0.89
OURS w/o MV attn 0.32 694.23 601.93 0.14 0.34 0.90 0.81
4D Gaussian (Wang et al., 2024a) 0.00 2102.52 2608.94 0.32 1.80 0.78 0.22
SVD (Blattmann et al., 2023) 605.10 648.41
SVD w/ MV attn 1151.05 931.33
Table 1: Multi-view 4D Video Generation Results. We compare our method with baselines in terms of cross-view consistency, RGB video generation quality, and depth generation quality. Our method consistently enables high-quality video and depth generation while maintaining strong cross-view consistency on both simulated and real-world datasets.
Figure 3: Qualitative Results and Comparisons under Novel Camera Views. Our method generates geometrically consistent 4D videos across camera views. In contrast, baseline results often exhibit significant cross-view inconsistencies or contain noticeable artifacts in the RGB or depth predictions. Video results can be found on the project website.
@@
Video prediction: To evaluate RGB video generation quality, we compute the commonly used Fréchet Video Distance (Unterthiner et al., 2018) (FVD) between the generated video and ground-truth video. FVD-$n$ evaluates the prediction from the reference view $v_{n}$, and FVD-$m$ evaluates the predcition from view $v_{m}$.
@@
Cross-View 3D Consistency: To evaluate the 3D consistency of the generated pointmaps across views, we compute the mean Intersection-over-Union (mIoU) on object masks. We use SAM2 (Ravi et al., 2024) to track the robot gripper and obtain binary masks from the generated videos for both views. For each frame, we lift the gripper mask in view $v_{n}$ to 3D space and then re-project it to view $v_{m}$. We then compute the IoU between the bounding boxes of the projected and original gripper masks. The mIoU is averaged over all time steps, and higher values indicate stronger 3D alignment across views.
Baselines. We compare our method with prior 4D generation approaches and variants of our model to evaluate generation quality and multi-view consistency. All models are trained on the same multi-view RGB-D video dataset and tested on novel viewpoints sampled from 100 unseen views per task. More details can be found in Section˜A.2.2.
@@
  • 4D Gaussian (Wang et al., 2024a): A baseline method that predicts one single-view RGB video using a finetuned SVD model (Blattmann et al., 2023) on our dataset, and then use a 4D Gaussian method, Shape of Motion (Wang et al., 2024a), to reconstruct a dynamic 4D scene from the video.
  • SVD (Blattmann et al., 2023): Stable Video Diffusion is a state-of-the-art video generation model. We finetune it on our dataset to predict stereo RGB video sequences.
@@
Results on Simulation Tasks. Results are reported in Table˜1, and Figure˜3 compares the predictions generated by different methods for the StoreCerealBoxUnderShelf task. Qualitative results and comparisons for other tasks can be found in Section˜A.4. Our method mostly achieves the best results across all tasks. For multi-view RGB video generation, it produces lower FVD scores than all baselines in most cases, indicating high temporal coherence and visual fidelity. Additionally, our method achieves the best depth prediction and cross-view consistency scores, demonstrating that the geometry-consistent supervision effectively guides the model to generate spatially aligned and geometrically coherent videos.
@@
We show that multi-view cross-attention is a crucial design choice for helping the model learn 3D geometric correspondences across camera views. The variant without multi-view cross-attention OURS w/o MV attn exhibits significantly lower 3D consistency, as measured by the mIoU metric. The video and depth generation quality also degrades in both views. In particular, the model performs poorly on the AbsRel-$m$ and $\delta_{1}$-$m$ metrics, indicating that it fails to learn the transformation needed to generate pointmaps in view $v_{n}$’s camera coordinate frame from view $v_{m}$ without the support of cross-attention layers. As shown in Figure 3, the gripper pose of OURS w/o MV attn is inconsistent across the two RGB views, and the projected gripper mask significantly misaligns with the actual gripper mask, as shown in the last column.
@@
4D Gaussian performs worse in video and depth generation quality, as well as cross-view consistency, as shown in Table˜1. In Figure˜3, the generated RGB frames appear blurry from unseen viewpoints, and the predicted depth is inaccurate. This is because the model is optimized on one single generated RGB video with fixed camera view, making it difficult to synthesize novel views.
@@
For the other baselines, SVD and SVD w/ MV attn, depth is not predicted and no geometric supervision is applied. While multi-view cross-attention is added to SVD w/ MV attn to enable information transfer between camera views, it operates over RGB features rather than pointmaps (which naturally encode 3D structure). As a result, the generated videos lack 3D geometric consistency across views. As shown in Figure˜3, SVD w/ MV attn produces a noticeable gripper mismatch between the two views and the RGB generation quality is significantly worse than our method.
@@
Results on Real-world Tasks. By fine-tuning the simulation-trained model on a multi-task real-world dataset for around 15k steps, our method is able to generate high-fidelity multi-view RGB-D sequences for several real world manipulation tasks. As shown in Table˜1 and Figure˜4, the generation results accurately capture both visual appearance and depth over time and consistently outperform baselines.
Figure 4: Real World Multi-Task 4D Generation Results. By fine-tuning the model trained in simulation, our model can generate high-fidelity real-world RGB-D sequences from multiple views that accurately capture the robot gripper motion over time. Video results can be found on the project website.